### v2

# Do Current Language Models Support Code Intelligence for R Programming Language?

Abstract.
Recent advancements in developing Pre-trained Language Models for Code (Code-PLMs) have urged many areas of Software Engineering (SE) and brought breakthrough results for many SE tasks. Though these models have achieved the state-of-the-art performance for SE tasks for many popular programming languages, such as Java and Python, the Scientific Software and its related languages like R programming language have rarely benefited or even been evaluated with the Code-PLMs. Research has shown that R has many differences with other programming languages and requires specific techniques. In this study, we provide the first insights for code intelligence for R. For this purpose, we collect and open source an R dataset, and evaluate Code-PLMs for the two tasks of code summarization and method name prediction using several settings and strategies, including the differences in two R styles, Tidy-verse and Base R. Our results demonstrate that the studied models have experienced varying degrees of performance degradation when processing R programming language code, which is supported by human evaluation. Additionally, not all models show performance improvement in R-specific tasks even after multi-language fine-tuning. The dual syntax paradigms in R significantly impact the models’ performance, particularly in code summarization tasks. Furthermore, the project-specific context inherent in R codebases significantly impacts the performance when attempting cross-project training. Interestingly, even when Large Language Models like CodeLlama and StarCoder2 are used for code generation, the Pass@K ($K=1,5,10$) results lags signigicantly behind Python scores. Our research shows that R as a low resource language requires different techniques to collect a high quality data. Specifically separating the two R styles has a great impact on the results and the separate dataset could increase the performance of the models. Our research sheds light on the capabilities of Code-PLMs and opens new research directions for researchers and practitioners for developing code intelligence tools and techniques for R. With R’s widespread use and popularity, the results of our study can potentially benefit a large community of R developers, both in research and industry.
Empirical Studies, Code Summarization, Method Name Prediction, R Programming Language, Code Generation in R, R programming Styles (Tidy-verse and Base)
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: TOSEMjournalvolume: 1journalnumber: 1article: 1publicationmonth: 10ccs: Software and its engineeringccs: Computing methodologies Natural language processing
1. Introduction
The R programming language was first developed by statisticians Ross Ihaka and Robert Jetman at the University of Auckland, New Zealand, in the early 1990s as a personal project (Ihaka and Gentleman, 1996). R was originally designed as a statistical programming language and a statistical teaching tool by combining the syntax of S, which is widely used among statisticians, with the semantics of Scheme  (Ihaka and Gentleman, 1996). The use and impact of the R language are growing every year as the number of users and application scenarios increase  (Fox and Leanage, 2016). Version 1.0.0 of R appeared on the CRAN (The Comprehensive R Archive Network) [^1] https://cran.r-project.org/ in 2000, and by the time this paper is written (2024), the current version of R is 4.4.1 and there are $\sim 21,423$ available widely used packages on CRAN.
With the growing size of the community and applications of R, it is ranked 6th in the PYPL Popularity of Programming Language  [^2] https://pypl.github.io/PYPL.html and 17th in the popularity of programming languages  [^3] https://spectrum.ieee.org/top-programming-languages-2024. R’s extensive library of packages, ranging from data manipulation to advanced statistical modeling, makes it one of the best choices in teaching (Wickham, 2014; Fox and Leanage, 2016), IT Sector (Nordmann et al., 2022), scientific fields (Medeiros Mirra et al., 2023; Nabi et al., 2022; Ningthoujam et al., 2023; Abdollahi et al., 2022), and Finance (Traykov et al., 2018). Though R is studied from various Software Engineering perspectives (Zanella and Liu, 2020; Huang and Yang, 2022; Plakidas et al., 2017; Vidoni, 2022), developing automation tools to support R developers, such as comment generation, is still in its infancy, compared to other languages like Java or Python.
The development of pre-trained language models has revolutionized the field of software engineering and development (Niu et al., 2023). These models, trained on massive datasets using deep learning techniques, possess the ability to generate human-like text. The significance of pre-trained language models in software engineering and development cannot be overstated (Zhu and Pan, 2019). They facilitate code autocompletion and generation, enhancing developers’ productivity and accuracy.
However, even though the R language has unparalleled importance in related fields, very limited research on code intelligence in R has been shown. A conventional GPT-2 (Radford et al., 2018) model is used for the code completion task in R (Popov et al., 2021). Though a large number of pre-trained models have been successfully developed in recent years and have proven their effectiveness in various software engineering (SE) tasks (Liu et al., 2019; Devlin et al., 2019; Feng et al., 2020; Guo et al., 2020; Wang et al., 2021b), to the best of our knowledge, no model has attempted to extend their effectiveness to the R language. While progress has been made in code intelligence for R (Popov et al., 2021), there exists a gap in the literature, limiting the realization of its full potential.
More recently, Large Language Models (LLMs) have demonstrated remarkable capabilities across various software engineering tasks (Radford et al., 2018; Brown et al., 2020; Li et al., 2023; Roziere et al., 2023). However, recent studies indicate that these models exhibit inconsistent performance across different programming languages (Ouyang et al., 2024). The performance disparity in different language is particularly evident in the R programming language, where the unique characteristics of the language pose specific challenges for LLMs. The coexistence of multiple syntax paradigms in R, i.e., Tidy-verse and Base, along with its heavy reliance on project-specific contexts and specialized libraries, could create additional complexities that current LLMs struggle to handle effectively.
In this study, we intend to explore to what extent the current techniques are applicable to R. However, as there is no current dataset, we first collect a dataset of high-quality R packages from GitHub repositories (inclusion/exclusion criteria described in Section 3.2.2). Then, we consider four pre-trained models, CodeBERT (Feng et al., 2020), GraphCodeBERT (Guo et al., 2020), UniXcoder (Guo et al., 2022) and CodeT5 (Wang et al., 2021b) to empirically study their performance for two tasks of code summarization and method name prediction, with various programming styles and settings. The results of the study are compared across models and with other programming languages. We then investigate an R-specific approach based on different programming styles to evaluate its effect on the models’ performance. In this study we also investigate the performance of LLMs for R language on code generation. This paper is the second Stage of our accepted TOSEM Registered Paper [^4] https://dl.acm.org/journal/tosem/registered-papers, in which we are reporting the results along with the discussions.
Our results reveal several key findings regarding the models and their capabilities for R programming language. First, studied models consistently experience varying degrees of performance degradation when processing R code compared to other programming languages. Second, our experiments demonstrate that multi-language fine-tuning does not uniformly improve performance across all PLMs for R-specific tasks, contrary to expectations. Third, we find that different syntax of the R styles in Tidy-verse and Base code significantly impact the performance, which is affecting the results more for code summarization task. The challenges extend beyond syntax to project-specific contexts inherent in R codebases, which poses a substantial challenge when attempting cross-project training. This is further complicated by the observation that combined syntax approaches generally underperform compared to either individual syntax paradigms. Our experiments with LLMs also show the low Pass@K scores for code generation, falling much behind the scores obtained for Python. These insights emphasize the need for more nuanced approaches to improving model performance for R programming, beyond simply increasing training data or applying multi-language fine-tuning strategies.
The contributions are as follows:
  • We open source an R dataset to support other research related to code intelligence in R [^5] https://zenodo.org/records/13871742.
  • We empirically evaluate the performance of different pre-trained models for R, for the specified tasks.
  • We provide an R-specific approach and study its effect on the models’ performance.
It is worth noting that though the foundation models (i.e., Large Language Models) in the past few months have found their way into SE, there is rarely a study or application available to use them for R. Additionally, their usage is costly if one wants to call their APIs. Therefore, their usage is not available for all, and fine-tuning them is not possible due to being large, besides their hidden ability for this programming language. LLMs have some shortcomings (Miceli-Barone et al., 2023) and still studying the smaller Code-PLMs can benefit the community. Therefore, we have studied the Code-PLMs and also added results for code generation using LLMs.
Significance: This study provides valuable insights into the effectiveness of using pre-trained code models for R and highlights the need for further research in this field. The results of our work can be used as a reference for future research and provide a foundation for the development of intelligent tools for the R community. Applying software engineering practices such as version control, documentation, testing, and modularization can lead to more reliable, maintainable, and scalable R code (Vidoni, 2021). Additionally, scientific software developers consider functional correctness, performance, portability, and maintainability as the top important characteristics (Johanson and Hasselbring, 2018). This is of specific importance as research shows that many scientists develop/use scientific software and “better software leads to better research” (Goble, 2014). Our work, thus, can provide insights to support the R developers, superficially in the context of program comprehension and updating comments, which could lead to better maintainability. This is one of the two important aspects for scientific software developers, and R is considered as a language used for the scientific software development (Arvanitou et al., 2021). We provide discussions and implication at the end, that helps shaping the future research for code intelligence in R, including the need to curate R datasets and developing separate techniques to increase the performance of the models, including LLMs, for the two R styles.
The rest of this paper is as follows. In Section 2, we provide an overview of the related research. We cover the methodology in Section 3. The result of our study is shown in Section 4. We discuss the results in Section 5 and the implications and threats to validity in Sections 6 and 7, respectively. Finally, we conclude this study in Section 8.
2. Related Works
In this section, we provide some background information and related work to our study.
2.1. Pre-trained Language Model for Code
In recent years, Pre-trained Language Models (PLMs) that are pre-trained on code –we refer to these models as Code-PLMs– have achieved great performance in different software engineering tasks (Wang et al., 2021b; Feng et al., 2020; Guo et al., 2020, 2022). PLMs are self-supervised deep learning models that have been previously trained on large amounts of text data to perform various natural language processing tasks. These models capture knowledge during the pre-training phase, which is then transferred to different downstream tasks, mainly by fine-tuning the model on specific downstream tasks.
Pre-trained models, specifically transformer-based models (Vaswani et al., 2017), have found significant applications in various software engineering tasks including code summarization, bug detection, and code translation (Feng et al., 2020; Guo et al., 2022). Examples of these models are CodeBERT (Feng et al., 2020), GraphCodeBERT  (Guo et al., 2020), CodeT5 (Wang et al., 2021b), and UniXCoder (Guo et al., 2022). PLMs outperform many of the state-of-the-art models on several datasets and tasks and have been used extensively in multiple software engineering tasks (Feng et al., 2020; Husain et al., 2019; Niu et al., 2023).
The use of pre-trained language models in software engineering offers several advantages, such as reducing the amount of labeled data required for training and increasing the efficiency of training (Han et al., 2021). They also provide a strong baseline for evaluating the performance of models and can be used to develop novel techniques in this field.
Due to the advantages of the PLMs for source code, there are works that investigate the usage and capabilities of Code-PLMs. Wan et al. (2022) studied the features captured by CodeBERT and GraphBERT through attention analysis, probing, and syntax tree induction, on Python, Java, and PHP. Troshin and Chirkova (2022) introduce probing tasks to discover the syntactic and semantic information of various Code-PLMs. Karmakar and Robbes (2021) uses probing tasks to study the Code-PLMs but does not include R language. Ahmed et al. (2023b) studies the Code-PLMs specifically by evaluating the models’ ability to learn semantics through objective and straightforward evaluation.
With all those studies on different areas and tasks in software engineering, to the best of our knowledge, no study has attempted to apply PLMs to the R programming language.
2.2. Scientific Programming Languages
Scientific programming languages play a pivotal role in the realm of data analysis, statistical modeling, and visualization, catering to the needs of researchers, statisticians, and data scientists (Staples, 2023). Among the diverse array of programming languages utilized in scientific domains, R stands out as a powerful and popular choice due to its robust statistical capabilities and extensive ecosystem (Giorgi et al., 2022). One of the key contributors to R’s widespread adoption is its vibrant and dedicated community, which has actively expanded the language’s functionality. Over the past decade, there has been a lot of research devoted to expanding the applicability and functionality of the R language. Notable contributions include the development of ggplot2 (Villanueva and Chen, 2019), a sophisticated and flexible data visualization package that enables the creation of high-quality graphics with ease. The Tidy-verse (Wickham et al., 2019), a collection of R packages designed to enhance data manipulation and analysis, has further streamlined workflows, promoting code readability and efficiency. RStudio, an integrated development environment (IDE) specifically tailored for R, has played a pivotal role in facilitating a user-friendly programming experience. Offering features like syntax highlighting, code completion, and interactive data visualization. Benefiting from the fact that these contributions have made the R language more applicable, many researchers from other fields have incorporated the R language into their analyses (de Jong et al., 2022; Zhao et al., 2021). Staples (2023) studied how R evolved, and showed that the emergence of the Tidy-verse is one of the primary causes behind the evolution of R. At the time of writing this manuscript, there is no research on the applicability and commonality of pre-trained models and large language models over R language.
2.3. Scientific Software, Jupyter Notebooks, and R Studies
In software engineering, several attempts and perspectives have been aimed at advancing code intelligence in R. However, to the best of our knowledge, no research has been conducted explicitly exploring the potential of pre-trained language models in this context.
Popov et al. (2021) develop a deep learning model for code completion in R. Sharma et al. (2022b) studied Self-Admitted Technical Debt (SATD) in R software packages and Chandramouli et al. (2022) presented analyzeR, a SonarQube plugin for analyzing object-oriented R packages. Zanella and Liu (2020) study the R packages on CRAN through a social network perspective. They report that the performance of R packages can be explained as a flow of information in the obtained network. Huang and Yang (2022) conducts an empirical study on the R ecosystem and found that the R development has been done by also software engineers and from other disciplines such as biology, environmental and medical science, other than statistics.
In another work, Plakidas et al. (2017) investigates the evolution of R and Ooms (2013) discusses the dependency versioning of R and demonstrates the issue and the solution through three use cases. Morandat et al. (2012) applies static and dynamic analysis to assess the design of R. Korkmaz et al. (2018) studies how the impact of R packages could be modeled through their dependencies and their contributors’ network.
Package dependency in R and CRAN packages (Decan et al., 2016), distribution of R packages (Decan et al., 2015), characterizing the bugs in R and Python (Ahmed et al., 2023a), and Studying the API breaking changes in a specific field of R packages (Chowdhury, 2023) are among other research on R. Babu (2022) develops a question answering system for R and (Wrenn et al., 2023) conducts the dependently typing for R vectors, arrays, and matrices.
Another closely researched area is scientific software and scientific computing. Ivie and Thain (2018) discussed the reproducibility in scientific computing and Goble (2014) reports that 91% of scientists surveyed online use scientific software, and at least 38% spend one fifth of their time developing scientific software. The author discusses that better software leads to better research.
Other works also study the data science practices in Jupyter notebooks. Liu et al. (2021) develop a model to generate code documentation for Jupyter Notebooks. Zhang et al. (2022) collect a dataset and build a model to classify the cells of Jupyter notebooks in data science projects. Yang et al. (2021) develop a Jupyter notebook extension based on program synthesis and and test selection to generate documentations that help users find bugs for data wrangling. Dong et al. (2021) study the data cleaning practices in 30 Jupyter notebooks. Robinson et al. (2022) examine Jupyter notebooks to define error identification strategies of Python notebooks. Their work is a replication of a similar work with user study to find these strategies for R Markdown notebooks Abdel-Ghani et al. (2022). Maintenance (Jiang et al., 2022), bug analysis (de Santana et al., 2022), structure improvement (Titov et al., 2022), and understanding documentation practices Wang et al. (2021a) are among other research explored in Jupyter notebooks.
Differences among our work and the current literature. Among the related works, the closest ones are the approaches that use deep learning to detect SATD (Sharma et al., 2022b) or complete code in R (Popov et al., 2021). The work multiple_E (Cassano et al., 2022) purposed a translater that is able to translate HumanEval (Chen et al., 2021) benchmark dataset to other programming language including R. Still, none of these or other research studies the ability of Code-PLMs for different tasks in R or considers the two R styles. Though the Jupyter notebook supports R programming language, the current studies mainly focus on Python-based libraries, excluding R. The other research on R programming language are related to empirical studies to gain a better insight about its different aspects. Our work differs with all the current literature, by first, providing a dataset for code summarization and method name prediction, and then exploring the capabilities of Code-PLMs for R programming language, in various settings including intra- and cross-projects and considering its different styles, Tidy-verse and Base.
3. Methodology
In this section, we will review the research questions, data collection process, studied tasks and Code-PLMs, along with our approach to answer each of the research questions.
3.1. Research Questions
The research questions we investigate and the answers in this research are as follows.
RQ1: What is the performance of the existing Code-PLMs on R language?
Many studies show that Code-PLMs obtain state-of-the-art performance for programming languages on the CodeSearchNet dataset. However, to what extent these models can accomplish the tasks for R language is not defined. Therefore, in this RQ, we evaluate the performance of the Code-PLMs for R. The answer to this RQ reveals the capability of the code-PLMs for R language, which is not explored in the existing research.
RQ2: Does the multilingual training in the fine-tuning phase increase the performance of the models compared to monolingual training for R?
Previous research (Ahmed and Devanbu, 2022) has shown that multilingual training in the fine-tuning phase increases the performance of the Code-PLMs for code summarization and method name prediction. In this RQ, we investigate whether such an effect is observed in the case of R. Answering RQ2 helps identify the effect of multilingual training for other languages beyond the six languages covered in the CodeSearchNet dataset (R in our study). This finding adds value to the current literature to assess the effectiveness of different approaches for low-resource languages, in this case, R.
RQ3: What is the effect of intra- or cross-project training?
Previous research (LeClair and McMillan, 2019; Ahmed and Devanbu, 2023) studied the effect of different dataset designs on the performance of the code summarization models. In this RQ, we investigate the outcome of models based on choosing the training and test functions from the same or different R repositories, also known as intra- or cross-project data selection. Although intra- or cross-project effect has been explored for other languages, it has not been explored for R. The results can shed light on designing experiments and the dataset to increase the performance of the models.
RQ4: Can we improve the performance of the models based on different programming styles in R?
Two of the popular programming styles in R are Tidy-verse and base R. The Tidy-verse is a collection of R packages designed for data science  [^6] https://www.tidyverse.org/. Though both are used frequently, the syntax of each one is different from the other (Wickham, 2014). Here, we investigate the effect of different programming styles on the performance of Code-PLMs. The results can benefit developing models for code intelligence tasks for the R developers, for both base and Tidy-verse.
RQ5: What is the performance of Large Language Models on the selected tasks for R?
Recent advances in foundation models or Large Language Models have shown promising results for some of the software engineering tasks (Brown et al., 2020; Roziere et al., 2023; Wang et al., 2021b) Accordingly, in this RQ, we will explore the performance of LLMs for the studied tasks for R. Currently, there is limited research on assessing the capabilities of LLMs for low-resource languages, including R. With the interest in using LLMs, the results of RQ5 will shed light on the potential of these models for code intelligence in R, compared to the code-PLMs.
3.2. Data Collection
3.2.1. Overview
Figure 1 presents an overview of our methodology for preparing the data. First, we collect GitHub repositories that include R files, considering the inclusion/exclusion criteria that we discuss next. Then, the eligible code snippets are parsed into code tokens using the tree-sitter parser. In the end, we followed the Roxygen2 documentation structure to match the code snippets and their corresponding natural language description to generate a dataset where each line of the data contains a pair of Programming Language (PL) and its correspondence Natural Language (NL).
Figure 1. Data collection workflow
3.2.2. Scraping GitHub Repositories
The API we used to collect the data is the GitHub REST API [^7] https://docs.github.com/en/rest?apiVersion=2022-11-28, which allows users to search for repositories based on specific criteria. Based on previous work (de F. Farias et al., 2016; Husain et al., 2019), we devised a similar selection strategy for the R language to ensure that the included code repositories are active, original, and relevant. To ensure the objectivity of the selection, we list the following Inclusion/Exclusion criteria adopted from previous works (de F. Farias et al., 2016; Husain et al., 2019).
Inclusion criteria:
  • The project should be a public project with an open-source license on GitHub.
  • The project needs to be an R package and have a proper R package structure. For example, it needs to contain a DESCRIPTION file.
  • The project needs to be an active and up-to-date project. We flirted the repositories where the last commit was made within a year.
Exclusion criteria:
  • The project should be a non-personal project. So all personal projects are excluded.
  • The project should not be archived. When a repository is archived, its issues, pull requests, code, labels and other components of the project become read-only. The owner and collaborators also can not update the project. So all archived repositories are excluded.
3.2.3. Data
To create the dataset for software engineering tasks, we formatted our data following the CodeSearchNet dataset (Husain et al., 2019) structure, where each line of the JSON file contains one data example, including the function name, the code tokens, the natural language description present in their respective documentation and the doc Strings. Each data point also contains metadata on the function, such as the repository it was extracted from. In order to do so, we performed a series of preprocessing and filtering on the raw data. The following discusses preprocessing and filtering in detail.
A total of $1,664$ GitHub repositories were collected using the GitHub API following the inclusion and exclusion criteria discussed above. We then pre-processed the data to obtain the required input for the software engineering tasks. Roxygen2888https://roxygen2.r-lib.org/, a popular documentation system for R that is similar to JavaDoc, was used to parse the functions into Abstract Syntax Trees (ASTs) and comments. We first make sure that these repositories are written in R and have a Roxygen framework. The Roxygen framework is needed to ensure the code comments could be extracted. For this purpose, we first inspect the file types of the source codes in our collected data. If the package includes .R files, typically found in an R folder in the root directory, it indicates the use of R language. Then, we look for Roxygen2 comments within these .R files. Roxygen2 is widely used and it manifests in the source code as specific comments, starting with the ‘#’ symbol, followed by the ‘@’ symbol, like ‘# @param’ or ‘# @return’. If we observe such comments, it suggests that the package employs Roxygen2 for documentation. Lastly, we examine the DESCRIPTION file in the package’s root directory. This file contains various metadata about the package, such as version number, maintainer, description, and package dependencies. If Roxygen2 is listed in the Imports or Suggests field, it provides a further indication that the R package employs Roxygen2. This process allowed us to extract the necessary information from the R code to form the input for the Pre-trained-models.
Programming language Training Dev Test
Ruby 24,927 1,400 1,261
JavaScript 58,025 3,885 3,291
Java 164,923 5,183 10,955
Go 167,288 7,325 8,122
PHP 241,241 12,982 14,014
Python 251,820 13,914 14,918
RBase 8,291 1,031 1,031
RTidy 24,380 3,047 3,048
RCombine 32,671 4,078 4,084
Table 1. CodeXGLUE and R dataset statistics
3.2.4. Parser
Roxygen is a documentation generator for R code that uses a parser to analyze R code and extract information that can be used to create documentation (Vidoni, 2022). Roxygen can use the information in the function to generate documentation for the function. For example, it can extract the names and default values of the function’s arguments, as well as information about the return value of the function. Roxygen can also use ASTs to identify any errors or potential problems in the code, such as syntax errors or unused variables.
In this study, we utilize Roxygen to achieve two primary objectives. First, we employ Roxygen to match code snippets from R language packages with their corresponding natural language descriptions. Second, we leverage Roxygen and the dependency information present in R language packages to classify them into either the base R or the Tidy-verse version. The process of using Roxygen to match code snippets with their corresponding natural language descriptions can be outlined as follows. Within R language packages that adhere to the Roxygen framework, there exist two directories: R and Rd (R document). For each R file, we specifically identify the method names within the code and search for matching method names in the Rd directory. Upon finding a match between a function name in the code and its counterpart in an Rd file, we establish a linkage. This enables us to obtain a comprehensive mapping between the code tokens within a function and the corresponding natural language description associated with that function.
Furthermore, we utilize Roxygen and Tree-sitter [^9] https://tree-sitter.github.io/tree-sitter/– GitHub’s universal parser– to classify an R language package as belonging to either base R or the tidy-verse. As previously described, the tidy-verse is a collection of R language packages that provide additional functionality and convenience on top of the R language itself. In R packages that potentially utilize the tidy-verse, their dependency environment includes the importation of R packages contained within the tidy-verse collection. To determine the classification, we employ Tree-sitter to parse and analyze the structure of the description file within an R language package. Specifically, we focus on the import section, which contains the necessary information. If this section includes the importation of the tidy-verse collection, we classify the R language package as belonging to the tidy-verse. Conversely, if the import section does not include the tidy-verse, the package is classified as base R.
After this processing, the dataset was reduced to 1583 repositories, pertaining to 40,833 functions (i.e., code-comment pairs). The statistics of our collected dataset are in Table 1. The Base R dataset is shown with RBase, the RTidy represents the Tidy-verse data, and the RCombine is the original collected data, without separating Base and Tidy-verse. The other rows of the table demonstrate the dataset from CodeSearchNet. Follwoing the previous works (Husain et al., 2019), we split the RBase, RTidy, and RCombine with 8:1:1, as shown in Table 1. Note that we have ensured that there are no duplicate functions in the dataset.
3.3. Tasks
In the following, we explain the selected tasks for our experiments.
Code Summarization is a technique in software engineering that aims to automatically generate a concise and informative summary of the code  (Feng et al., 2020; Guo et al., 2020). Code summarization has the advantage of being fast, consistent, and reproducible, and can help developers easily understand the semantics of the source code (Zhu and Pan, 2019). This can be particularly useful in large and complex software projects where it can be difficult to keep track of the different components and understand how they interact with each other. In recent years, code summarization has gained significant attention in the software engineering community, with several research studies exploring the effectiveness of various summarization techniques and models (Devlin et al., 2019; Husain et al., 2019). However, previous research in code summarization has mainly focused on programming languages that are widely used in firms such as Java and Python  (Feng et al., 2020; Guo et al., 2020; Liu et al., 2019), leaving a gap in our understanding of code summarization for other programming languages, such as R. Therefore, we study the code summarization as one of the tasks. The dataset we collected includes both the methods and their natural language comments, written by developers. Therefore, we can use the collected dataset and it is ready to train the models for code summarization.
Method Name Prediction or extreme code summarization refers to the task of predicting or generating appropriate names for methods (functions or procedures) in programming code (Allamanis et al., 2016). Given a code snippet, the model aims to generate a suitable name that accurately represents the intended behavior or functionality of the method. This task is important in software development as well as code comprehension because well-named methods can improve code readability, maintainability, and understandability (Allamanis et al., 2016). Method Name Prediction often involves training models on large-scale code repositories to learn patterns and common naming conventions used in programming. The models can then use this knowledge to generate plausible method names for given code contexts. This task has applications in various code-related tasks, such as code completion, code summarization, and code search, where generating informative method names can greatly assist developers in writing or understanding code more efficiently (Petukhov et al., 2022). It also has been studied in multiple previous research (Ahmed and Devanbu, 2022; Alon et al., 2019, 2018). Therefore, we chose this task in our study. The collected dataset is valid for the method name prediction task, as given the code without the method name, and the natural language comments, we can train and evaluate the models to generate the method name.
Code Generation aims to generate a code snippet from a given natural language description. Ridnik et al. (2024) purposed AlphaCodium which performed well on code generation benchmark. The model is given as input a description written in natural language, and the goal is to generate the code that corresponds to the description. Developers often take code from blogs, posts, articles, and other websites and adapt it to their context. The code task generation task is important as it provides a better replacement solution by generating these codes for users based on natural language descriptions. We select code generation to assess the capabilities of the models for a different category of code intelligence task (natural language to Code task). As there is no code generation for R, we will follow the HumanEval (Chen et al., 2021) dataset, which is in Python. We use MultiPL-E (Cassano et al., 2023) to evaluate the LLM’s performance on R. MultiPL-E (Cassano et al., 2023) is a multi-language benchmark that translates the HumanEval dataset from Python into 18 other programming languages. We sample 20 completions per prompt at temperature 0.2 with top-p 0.9 following (Cassano et al., 2023). Note that typically 50 is needed to compute Pass@1 but Cassano et al. (2023) reports that 20 completions would give a stable estimate for the Pass@1 score. Note that due to a lack of training data for code generation in R, we will run the experiments in a zero-shot paradigm, meaning that the models will not be fine-tuned for this task. Instead, we only ask the model to generate code without showing an example (i.e., zero-shot).
3.4. Studied Code-PLMs
In this section, we will briefly discuss the chosen benchmarking pre-trained models. We choose these models based on their architecture, pre-training tasks, and performance, and we follow the recommendation of a recent empirical study for selecting them (Niu et al., 2023). Niu et al. (2023) conducted empirical studies on various code intelligence tasks and code-PLMs and provided recommendations for the model choices for each task. Accordingly, we choose CodeBERT, GraphCodeBERT, UniXcoder, and CodeT5 in our study. Note that at the time of this study, to the best of our knowledge, there is no research on evaluating the models for R language. Therefore, we follow the recommendations of (Niu et al., 2023) and choose the following models, in addition to conducting experiments with large language models, GPT-3.5-Turbo and CodeLlama.
CodeBERT (Feng et al., 2020) is a language model specifically designed for programming languages. CodeBERT’s architecture is based on the Transformer model, similar to the original GPT model, but with modifications to accommodate programming code. CodeBERT is pre-trained using two primary tasks: masked language modeling (MLM) and replaced token detection (RTD). For MLM, which has been shown to be effective in literature (Devlin et al., 2019), random tokens in the input code are masked, and the model is trained to predict the masked tokens. For RTD (Clark et al., 2020), the objective is to use both bimodal and unimodal data for training. CodeBERT has shown promising results in various programming language-related tasks, such as natural language code retrieval, and code summarization. It has achieved state-of-the-art performance at the time of publication in code-related benchmarks and competitions, demonstrating its effectiveness in understanding and generating code.
GraphCodeBERT (Guo et al., 2020) extends CodeBERT (Feng et al., 2020) by incorporating the graph-based representation of code. It represents code snippets as abstract syntax trees or control flow graphs (CFGs) to capture the structural information of the code. GraphCodeBERT follows BERT(Devlin et al., 2019) and uses multi-layer bidirectional Transformers as the backbone. In addition to source code, GraphCodeBERT incorporates paired comments during pre-training to enhance the model’s ability to handle code-related tasks that involve natural language, such as natural language code search (Feng et al., 2020). Moreover, it includes data flow, represented as a graph, as an input component for the model. GraphCodeBERT employs similar pre-training tasks as CodeBERT. It also incorporates Edge Prediction and Node Alignment to enhance the model’s understanding of relationships between entities mentioned in the text. By pre-training the model on a task that involves predicting edges between entities or concepts, the model can potentially learn to capture the semantic connections and dependencies between them. GraphCodeBERT has demonstrated improved performance over CodeBERT (Feng et al., 2020) in tasks that require a deeper understanding of the code structure. It has been successful in tasks such as code translation, code clone detection, and code refinement.
UniXcoder (Guo et al., 2022) is a model designed for multilingual code-mixing scenarios, where programming code contains multiple languages. It aims to handle code written in multiple programming languages within the same codebase. UniXcoder employs a multi-modal architecture that combines the strengths of both Transformer models and recurrent neural networks (RNNs). It consists of a hybrid Transformer-RNN model, where the Transformer processes the textual code tokens, and the RNN handles the language-related features. UniXcoder utilizes pre-training tasks similar to CodeBERT to learn code-related representations. Additionally, it incorporates tasks specific to multilingual code-mixing scenarios, such as language identification and code-switching prediction. UniXcoder has demonstrated effectiveness in handling multilingual code-mixing scenarios and achieving high performance in code-related tasks involving multiple programming languages.
CodeT5 (Wang et al., 2021b) is another variant of the T5 (Text-To-Text Transfer Transformer) model specifically designed for code-related tasks. It combines the power of T5’s text-to-text transfer learning with code-specific pre-training. CodeT5 utilizes the T5 Transformer architecture (Wang et al., 2021b), which consists of a transformer encoder-decoder framework. The model takes code as input and generates code-based outputs for various downstream tasks. CodeT5 (Wang et al., 2021b) employs a text-to-text transfer learning paradigm, where a variety of code-related tasks are cast into a unified text generation format. These tasks include code translation, code summarization, code completion, and more. The model is trained to generate the target code given the corresponding input and task description. CodeT5 (Wang et al., 2021b) has achieved strong performance in multiple code-related tasks, surpassing previous state-of-the-art models. CodeT5 (Wang et al., 2021b) has shown effectiveness in code generation, code summarization, and code translation tasks across various programming languages.
CodeLlama-7B. CodeLlama-7B is a model from the Code Llama (Roziere et al., 2023) model series which is a large programming domain-oriented language model based on Llama 2 (Touvron et al., 2023), which allows for the direct generation or understanding of code using textual hints (Prompts). CodeLlama has the ability to perform code completion by generating up to 100k tokens (Roziere et al., 2023). Additionally, CodeLlama has the ability to generate zero-sample instructions for programming tasks. In addition, CodeLlama has zero-sample instruction adherence for programming tasks including code generation. This small-sample learning capability is worth noting when testing code to accomplish tasks in small resource languages like R. Roziere et al. (2023) claim that CodeLlama can be used for a wide range of programming tasks and is one of the best performing LLM at public programming tasks, making developers’ workflows faster and more efficient, and lowering the learning threshold for programming. Therefore, we select CodeLlama to explore the RQ5 answer.
StarCoder2-base. In recent years, the field of code generation has advanced significantly with the development of Large Language Models (LLM) tailored for code-related tasks. One of the most prominent models is StarCoder (Li et al., 2023), along with its base variant, StarCoderBase, both part of the BigCode project [^10] https://huggingface.co/bigcode.StarCoder and StarCoder-base are models trained on permissively licensed data from GitHub, covering over 80 programming languages, Git commits, GitHub issues, and Jupyter notebooks. StarCoder-Base, which has approximately 15 billion parameters, was trained on 1 trillion tokens. StarCoder-Base was further fine-tuned on 35 billion Python tokens to create StarCoder, specifically optimized for Python-related tasks.
StarCoderBase surpasses many existing open-source Code LLMs on popular programming benchmarks and competes with closed models like OpenAI’s code-cushman-001 (the Codex model behind early versions of GitHub Copilot). With a context length exceeding 8,000 tokens, StarCoder can handle large inputs, making it ideal for tasks like acting as a technical assistant, autocompleting code, modifying code based on instructions, or providing natural language explanations of code snippets. Given Starcoder’s performance and versatility, StarCoder models offer a strong foundation for community adoption and customization across various applications including code generation. Therefore, we use StarCoder-base in our experiments in RQ5.
Codegen-2B-multi. We should note that, initially we intended to use CahtGPT as one of the LLMs for code generation. Since GPT-3.5 (chatgpt) is not open-source, we used CodeGen-2B-multi as an alternative for the code generation task. CodeGen-2B-multi is a large-scale transformer model trained on both natural language and multiple programming languages, allowing it to generate code from natural language descriptions. CodeGen-2B-multi is open-source, making it a valuable resource for research and development. The model is pre-trained on natural language datasets and fine-tuned on multilingual code datasets, enabling it to perform well across various programming languages for tasks like code completion and synthesis. Therefore, we used it for the experiments in RQ5.
3.5. Approach
In this section, we will provide a detailed experimental setup of each of the studied RQs.
RQ1: This RQ is motivated by the lack of research and understanding of the performance of Code-PLMs for the R programming language. To address this challenge, we compared the four Code-PLMs for each of the languages on the CodeXGLUE benchmark (which is the same data as CodeSearchNet for these tasks), and the R dataset. To achieve the comparison, we trained each of the models on each of the programming languages shown in the first six rows of Table 1 and also on the RCombine datasets separately, and then tested on the test split of each one. Therefore, we can evaluate them when trained and tested on each of the languages in a mono-lingual setting.
RQ2: For this RQ, we intend to evaluate the effect of multi-lingual fine-tuning for R. To achieve this, we combined the training splits of all the seven languages (i.e., Ruby, JavaScript, Java, Go, PHP, Python, and RCombine) and then train each of the Code-PLMs on this multi-lingual dataset. We tested the performance of each model on the test set of each of the programming languages separately. The training and testing were done for code summarization and method name prediction, and the results are reported accordingly.
RQ3: For this RQ, we use the RCombine dataset and we split the functions of the training and test set based on the repository they belong to. For the intra-project setting, we spited the training and test sets such that we have some functions from every repository in the training set and some of its functions in the test set. Note that we take care of not having duplicate functions in the training and test sets. For the cross-project setting, the repositories are split such that their functions only belong to the training or test set. Therefore, the model does not have functions from the same repository in the training and test set.
RQ4: As mentioned earlier in section 3, there are two different opinionated programming styles in R, the Base version, and Tidy-verse. The Base R is the fundamental R programming language, and it is the only dialect for a long time it can be more stable when compared with Tidy-verse (Wickham, 2014). Tidy-verse, on the other hand, is a collection of packages that work as an add-on to base R that provides extra functionality and makes the workflow and syntax more user-friendly. Despite the benefits and user groups of each, the syntax can be very different. Listing 1 provides an example of Tidy-verse and a Base R example is shown in Listing 2. Based on these differences, we hypothesize that separating the datasets might affect the obtained scores in a positive way.
Therefore, for this RQ, we repeat the experiments of RQ1–RQ3, but with RBase and RTidy datasets separately. To have a fair comparison, we decrease the size of the RTidy to include the same number of records as the RBase. This will be in addition to having the experiments on the RTidy dataset with its original number of records.
RQ5: To answer this research question, we run experiments using CodeT5 (Wang et al., 2021b) and UniXCoder (Guo et al., 2022), as well as ChatGPT (Brown et al., 2020), CodeLlama (Roziere et al., 2023) and StarCoder2 (Li et al., 2023). We exclude other models as they are not developed for code generation. For code generation, as stated earlier, the models will be prompted to generate code. We prompted ChatGPT (Brown et al., 2020) and CodeLlama in a zero-shot setting and compare its results with the full fine-tuned models. For code summarization and method name prediction tasks, we selected 100 samples from the test dataset (same as the ones used for human evaluation, described in Section 3.7).
Listing 1: Tidy-verse
1mtcars %>%
2 group_by(cyl, gear) %>%
3 summarise(mpg.mean = mean(mpg),
4 mpg.sd = sd(mpg),
5 wt.mean = mean(wt),
6 wt.sd = sd(wt)) %>%
7 ungroup()
Listing 2: Base R
1mtcars_by <- by(mtcars,
2 INDICES = list(mtcars\$cyl, mtcars\$gear),
3 FUN = function(x){
4 data.frame(cyl = unique(x\$cyl),
5 gear = unique(x\$gear),
6 mpg.mean = mean(x\$mpg),
7 mpg.sd = sd(x\$mpg),
8 wt.mean = mean(x\$wt),
9 wt.sd = sd(x\$wt))
10 })
11
12do.call(rbind, mtcars_by)
3.6. Evaluation Metrics
For code summarization, following previous works (Feng et al., 2020; Wang et al., 2021b), we evaluate the results using smoothed BLEU-4 score (Lin and Och, 2004). The smoothed BLEU-4 score is a technique used in the field of machine translation to assess the quality of translations. It is an extension of the BLEU (Bilingual Evaluation Understudy) metric, which measures the similarity between a machine-generated translation and one or more reference translations. The BLEU score can be calculated using the following formula1:
(1) $BLEU=BP*exp(\sum_{n=1}^{N}W_{n}logP_{n})$
where BP is the brevity penalty and $P_{n}$ is the modified precision score, let c be the length of the candidate and r be the effective reference corpus length, BP can be calculated using:
(2) $BP=\left\{\begin{array}[]{lr}1&\textsf{if}\ c>r\\ e^{1-r/c}&\textsf{if}\ c\leq r\\ \end{array}\right.$
When reporting the smoothed BLEU score, the model with the higher score is considered a better model for this task. A higher score is considered better.
For method name prediction, we follow the previous work (Ahmed and Devanbu, 2022), and report the F1-score to evaluate the classification models, as defined below.
(3) $F1=2*\frac{precision*recall}{precision+recall}$
Where $precision=\frac{Truepositive}{Truepositive+Falsepositive}$ and $recall=\frac{Truepositive}{Truepositive+Falsenegative}$. Here, True positive indicates that the token that is part of the name is been predicted as part of the function name. When reporting the F1 score, the model with the higher F1 score is considered a better model for this task.
For code generation, following the metrics used in code generation studies using HumanEval (Roziere et al., 2023), we report the pass@K. Pass@K is an evaluation metric used to measure the probability that at least one correct answer is included among the top k responses generated by a model, calculated as follows (Chen et al., 2021):
(4) $\operatorname{pass}@k:=\mathbb{E}_{\text{Problems }}\left[1-\frac{\left(\begin{array}[]{c}n-c\\ k\end{array}\right)}{\left(\begin{array}[]{c}n\\ k\end{array}\right)}\right]$
For instance, if $k$ is set to $5$, pass@5 signifies the probability that at least one correct answer is present within the first five responses generated by the model. This metric helps assess whether the model’s output contains the correct result among multiple possibilities, providing a measure of the reliability of the model’s output. We will apply the metric for $k\in\{1,5,10\}$.
Statistical tests. Following the previous works (Ahmed and Devanbu, 2022), we apply the Wilcoxon signed-rank test to compare the results of the models for all the tasks.
3.7. Human Evaluation
Other than automatic evaluation metrics, we also conduct a human evaluation, following the protocol used in previous studies (Shahbazi and Fard, 2023; Sharma et al., 2022a). We randomly select $100$ samples from the test dataset for each task. Each sample will include code, ground truth values, and the generated output (code comments or method names) for all models. Two Computer Science graduate students, who are both familiar with R, conducted the evaluation. Each record was ranked by two reviewers and the average value were reported. Each sample was evaluated based on the following criteria, on a Likert scale of $1$ to $5$, where $1$ shows the lowest and $5$ shows the highest.
For code summarization, we considered the following aspects:
  • Informativeness (I), which considers how well the generated comment covers the key points of the code.
  • Relevance (R), which evaluates the consistency of the provided details in the comments and compares it with the details in the code.
  • Fluency (F), which assesses English fluency in terms of the comments being well-written and grammatically correct.
For method name prediction, we considered the semantic similarity of the ground truth and the generated names. To avoid bias, we hided whether the text (comment or method name) is generated by a model or is written by a human. Moreover, the evaluators were educated before scoring. We provided examples and explanations for each criterion before they start evaluation.
3.8. Experimental Setup
All the experiments were run on the UBC Advanced Research Computing Sockeye and Digital Research Alliance of Canada. Each job requires the allocation of a compute node with 1 CPU core, 48GB of memory, and 2 NVIDIA Tesla V100 32GB GPUs with 32GB of memory for execution. For hyper-parameters used during the training phase, we followed the setting introduced in the original papers. We trained the models for 10 epochs with a learning rate of $0.00005$. The source_length is set to $256$ and the target_length is to $128$ for code summarization. For method name prediction, the target length is adjusted to $10$. The batch size is set to $32$ for CodeBERT (Feng et al., 2020) and GraphCodeBERT (Guo et al., 2020), $48$ for UnixCoder (Guo et al., 2022) and CodeT5 (Wang et al., 2021b) to perform a fair comparison to the original paper and the between R language and other programming languages. All experiments are run on Dell EMC R440 GPU compute nodes.
4. Results
The obtained results are presented in this section.
4.1. RQ1: Performance of Code-PLMs in Mono-Lingual Fine-tuning Setting on R
4.1.1. Code Summarization
Source CodeBERT GraphCodeBERT UniXCoder CodeT5
JS 14.67 14.84 15.44 16.17
Go 18.12 18.48 19.26 19.25
Java 18.65 19.24 20.14 20.53
PHP 25.36 25.75 26.52 26.13
Ruby 11.54 11.76 15.01 15.42
Python 18.29 18.31 19.17 19.32
Combinedsix 19.15 19.20 19.89 19.73
RCombine 11.80 11.02 18.43 15.4
RTidy 17.92 10.38 18.93 16.64
RBase 19.55 9.56 20.93 17.61
Table 2. Performance for all the baselines for the six programming languages on code summarization, R styles, and the combined R dataset under monolingual training. The ‘Combined’ represents the combined dataset of the six programming languages from CodeXGLUE. ‘Rcombine’ shows the results when the data from both styles of R are combined together. Best results in each row are bold. For the three R datasets, we also underlined the best results in each column (only among the R datasets).
Table 2 presents the BLEU-4 scores for code summarization, for all languages. The Combinedsix represents the results when we combined all six languages from CodeSearchNet. Row Rcombine shows the dataset that we combine both Tidy-verse and Base R datasets.
We first replicated the summarization task by fine-tuning the pre-trained model for six languages, following the reported set of instructions. This step is done to ensure the reliability of our results and avoid errors that might have been introduced by the hardware or the system environment. For all six languages and models, our replicated results are close to the models’ reported values. For all models, the same training setup is used for R.
As shown in Table 2, for all the six languages, JavaScript (JS), Go, Java, PHP, Ruby, and Python, CodeT5 has the best performance. For some of the languages, UniXCoder has close scores, all being higher than CodeBERT and GraphCodeBERT. When the models are tested on RCombine, we observe that UniXCoder has the highest performance, being $18.43$ followed by CodeT5 BLEU-4 score of $15.42$. The CodeBERT and GraphCodeBERT both have lower score in the range of $11$ BLEU-4 scores.
It is worth noting that GraphCodeBERT (Guo et al., 2020) is primarily designed to model the structural relationships in code (e.g., how different components of code like functions, variables, and classes are related). This is beneficial for tasks that rely heavily on understanding the syntactic and semantic structure of code, such as code completion or defect detection. For summarization, where the goal is to generate natural language descriptions from code, a model needs to excel in both code understanding and natural language generation. Though we tokenizaed the R dataset using the same tool, Tree-sitter, GraphCodeBERT continues to struggle for R. A possible reason could be that R, as a programming language, is fundamentally different from other languages. This difference means that R does not share the same expressions or patterns, making it challenging for the model to fit the data effectively during training. So this might be a culprit for low performance of GraphCodeBERT on R.
4.1.2. Method Name Prediction
Source CodeBERT GraphCodeBERT UniXCoder CodeT5
JS 0.37 0.39 0.34 0.34
Go 0.52 0.52 0.49 0.40
Java 0.51 0.52 0.45 0.42
PHP 0.50 0.54 0.42 0.47
Ruby 0.42 0.45 0.41 0.33
Python 0.44 0.46 0.43 0.40
Combinedsix 0.43 0.46 0.35 0.38
RCombine 0.25 0.32 0.26 0.29
RTidy 0.26 0.31 0.25 0.29
RBase 0.26 0.32 0.23 0.30
Table 3. F1-score of the models for the six programming languages, R styles, and combined R dataset on name prediction under monolingual fine-tuning. Best results in each row are bold and best results in each column are underlined.
We study the performance of fine-tuning all models using RTidy, RBase and Rcombine. As we did for code summarization, we re-trained the baseline models, to get performance numbers for each of the six programing language in the codeSearchNet (Husain et al., 2019) dataset and report all the results. Table 3 shows the F1-score of all the all the programming languages for method name prediction. GraphCodeBERT outperforms CodeBERT, UniXCoder and CodeT5 in all languages except Go which GraphCodeBERT and CodeBERT show same scores. The performance of GraphCodeBERT demonstrate that GraphCodeBERT’s graph-based approach is more effective than the other three base line models in handling method name prediction. On the other hand, Unixcoder does not continue to maintain its performance in code summarization, instead, UniXCoder shows its struggle for method name prediction, being the lowest in JS, PHP, RTidy, RBase and Rcombine. UniXcoder is designed to handle multiple code intelligence tasks in a unified manner, shows varied results across the languages. UniXcoder’s scores lower than the other models, is potentially due to its broader generalization approach, which may not specialize as effectively in code understanding. CodeT5, likewise, did not continue to maintain its performance in this task, with CodeT5 being the worst performer in the other four languages: Go, Java, Python and Ruby.
The R language datasets present significant challenges for all models. The scores are around $0.25$ for most models and around $0.32$ for the best model (i.e., GraphCodeBERT). Interestingly, for this task, there is no difference between the RTidy, RBase and Rcombine; and all R datasets have exact or very close scores. This indicated the need to handle R’s domain-specific syntax and statistical focus by further adaptation or specialized pre-training models that works for R.
4.2. RQ2: Performance of Code-PLMs in Multi-lingual Fine-Tuning Setting on R
4.2.1. Code Summarization
CodeBERT GraphCodeBERT UniXCoder CodeT5
Language BLEU p-value BLEU p-value BLEU p-value BLEU p-value
JS 15.1 $0.008^{*}$ 15.77 $<0.001^{*}$ 15.63 0.452 16.03 $0.036^{*}$
Go 18.38 0.672 19.15 $<0.001^{*}$ 18.31 $<0.001^{*}$ 6.16 $<0.001^{*}$
Java 18.73 $0.003^{*}$ 19.84 $<0.001^{*}$ 19.58 $<0.001^{*}$ 7.39 $<0.001^{*}$
PHP 25.34 0.192 26.35 $<0.001^{*}$ 25.97 $<0.001^{*}$ 7.39 $<0.001^{*}$
Ruby 14.78 $<0.001^{*}$ 14.98 $<0.001^{*}$ 14.31 $0.004^{*}$ 5.63 $<0.001^{*}$
Python 18.59 $<0.001^{*}$ 19.01 $<0.001^{*}$ 18.88 $<0.001^{*}$ 7.76 $<0.001^{*}$
RCombine 7.24 $<0.001^{*}$ 9.65 $<0.001^{*}$ 14.34 $<0.001^{*}$ 2.33 $<0.001^{*}$
RTidy 8.4 $<0.001^{*}$ 11.14 0.766 18.79 0.077 2.23 $<0.001^{*}$
RBase 9.02 $<0.001^{*}$ 12.02 $0.003^{*}$ 20.02 0.759 2.11 $<0.001^{*}$
Table 4. Performance of all the baselines for the six programming languages on code summarization, R styles, and the combined R dataset under multi-lingual fine-tuning. Best results in each row are bold and best results in each column are underlined.
Table 4 represents the results for code summarization under multi-lingual fine-tuning. In this setting, the performance of the baseline models contradicted with previous findings (Ahmed and Devanbu, 2022), where multi-lingual fine-tuning typically improved model performance. For UniXcoder, and CodeT5, multilingual fine-tuning actually leads to performance degradation. UniXcoder experience a drop in performance for most programming languages, and CodeT5 shows slightly performance degradation only in Ruby, with significant drop in all other languages, indicating that the effectiveness of multilingual fine-tuning varies depending on the model. This drop in performance is also reported in previous works (Wang et al., 2023), and dropped even further with the addition of R language to the dataset. Specifically, for CodeBERT, the BLEU-4 scores increased for JavaScript, Go, and Ruby by $7\%$, $1.4\%$, and $28\%$ respectfully, the improvement is significant for JavaScript and Ruby, showing a modest improvement across these languages. However, other languages exhibited mixed results. For example, PHP and Java received almost no change, and for R language, the combined R dataset (RCombine) demonstrated a $39\%$ performance drop compared to mono-lingual setting.
When examining GraphCodeBERT, there was a consistent trend of performance improvement across all languages except for combined R dataset (RCombine), which is decreased by $1.4$ BLEU-4 score. All six languages of CodeSearchNet data have about one BLEU score improvement, except Ruby which jumps to $14.98$ from $11.76$, seeing the highest increase. The performance on R datasets was more variable, with RBase seeing a notable increase of $25\%$ (p-value equal to 0.003), while RTidy only increases $7\%$ (with p-value equal to 0.766 which is not significant). Here we can see that although GraphCodeBERT performed well under multi-lingual fine-tuning, the combination of RTidy and RBase would hurt the effects of fine-tuning while testing them, as shown in the results of RCombine.
In multi-lingual setting, UniXCoder (Guo et al., 2022) on the other hand, has reduced the BLEU scores for all languages, dropping the scores slightly. Particularly for RCombine, while it had previously demonstrated strong performance in monolingual settings, its the performance dropped by 22% (with p-value less than 0.001). For RBase and RTidy, the performance dropped slightly. This observation again highlights the negative effects from combining RTidy and RBase datasets during testing.
Lastly, CodeT5 demonstrated very low performance on all languages, except JavaScript. This performance drop is more noticeable for R datasets, where the BLEU scores are around $2$. The low scores could be attributed to CodeT5’s text-to-code architecture, which may not generalize as well in multi-lingual fine-tuning. While CodeT5 performed well in the monolingual setting, its performance dipped significantly during fine-tuning on the R datasets. This points to a potential limitation in its ability to handle diverse programming paradigms across different languages. It is worth pointing out that CodeT5 also took the most time to fine-tune, and during training we observed that the model had a hard time converging to the point where the model showed a drop in performance on the dev set. Combined with the performance of other base line models, this is a good indication that pre-trained models such as CodeT5 have some shortcomings in learning a language such as R, which has its own unique syntax and structure, specially for RCombine.
To evaluate the significance of the differences in model performance, we employed the Wilcoxon signed-rank test, a non-parametric statistical test commonly used for comparing paired samples. Wilcoxon signed-rank test was chosen due to its suitability for comparing two related samples, making it ideal for performance comparisons across different models on the same dataset. The p-value is computed by applying the Wilcoxon signed test on mono-lingual fine-tuning result and multi-lingual fine-tuning for code summarization and method name prediction. Under this setting the p-value represents whether the observed differences in performance are statistically significant or not. To control for the risk of false positives due to multiple comparisons, we applied the Benjamini-Hochberg (B-H) correction following previous study (Ahmed and Devanbu, 2022). The B-H correction is a widely used method for controlling the false discovery rate (FDR), which refers to the expected proportion of incorrectly rejected null hypotheses (false positives) among all rejected hypotheses. By adjusting p-values obtained from the Wilcoxon signed-rank test, the B-H correction allows us to maintain a balance between discovering true effects and limiting the inclusion of false positives when performing multiple comparisons.
The p-values in Table 4 that are $<0.05$ and show the numbers are statistically significant from mono-lingual setting, are shown with an asterisk (*). In the table, some results are not significant based on their p-value. These include JavaScript for UniXCoder, Go for CodeBERT, PHP for CodeBERT, RTidy for GraphCodeBERT and UniXCoder, and RBase for UniXCoder, these higher p-values indicate that the differences in BLEU scores for these models are not statistically significant. Most results though, show a significant difference for R.
In summary, multi-lingual fine-tuning presented varied impacts across models and languages. While certain models like GraphCodeBERT excelled in handling multi-lingual, all models struggled with RCombine. This suggests the disadvantages of the base line model, particularly when dealing with diverse languages and paradigms like those in the R language.
4.2.2. Method Name Prediction
CodeBERT GraphCodeBERT UniXCoder CodeT5
Model F1-score p-value F1-score p-value F1-score p-value F1-score p-value
JS 0.13 $<0.001^{*}$ 0.15 $<0.001^{*}$ 0.16 $<0.001^{*}$ 0.32 $<0.001^{*}$
Go 0.13 $<0.001^{*}$ 0.13 $<0.001^{*}$ 0.14 $<0.001^{*}$ 0.12 $<0.001^{*}$
Java 0.15 $<0.001^{*}$ 0.16 $<0.001^{*}$ 0.16 $<0.001^{*}$ 0.115 $<0.001^{*}$
PHP 0.21 $<0.001^{*}$ 0.20 $<0.001^{*}$ 0.19 $<0.001^{*}$ 0.107 $<0.001^{*}$
Ruby 0.14 $<0.001^{*}$ 0.15 $<0.001^{*}$ 0.18 $<0.001^{*}$ 0.112 $<0.001^{*}$
Python 0.16 $<0.001^{*}$ 0.16 $<0.001^{*}$ 0.19 $<0.001^{*}$ 0.114 $<0.001^{*}$
RCombine 0.20 $<0.001^{*}$ 0.28 $<0.001^{*}$ 0.23 $<0.001^{*}$ 0.10 $<0.001^{*}$
RTidy 0.21 $<0.001^{*}$ 0.38 $<0.001^{*}$ 0.23 0.227 0.097 $<0.001^{*}$
RBase 0.21 $<0.001^{*}$ 0.41 $0.002^{*}$ 0.24 $0.008^{*}$ 0.095 $<0.001^{*}$
Table 5. Result of method name prediction for RQ2. Best results in each row are bold and best results in each column are underlined. The p-values compared to mono-lingual setting are shown for each model. P-values smaller than $0.05$ show a statistically significant difference and are marked with an asterisk (*).
Table 5 shows the result of all the baselines for the six programming languages, R styles, and combined R dataset on method name prediction under multilingual training. We observe significant drop in performance for all four models and the six languges of CodeSearchNet dataset (the p-value is less than 0.001 for nearly all of models and language). Multilingual training generally provides models with a broader understanding of diverse languages by exposing them to a combined dataset, which helps in cross-linguistic generalization but may dilute performance on individual languages due to the need to balance between multiple linguistic patterns. For all six languages of the CodeSearchNet dataset, and for all models, the performance has dropped to $0.10s$ with only a few close to $0.2$.
In examining the R styles, especially RTidy and RBase, the results under multilingual training are notably different. GraphCodeBERT performs surprisingly well in these cases, boosting the performance; particularly under multilingual training, where it scores $0.38$ for RTidy and $0.41$ for RBase, compared to $0.31$ and $0.32$ in mono-lingual setting, which is statistically significant. This means the performance is improved by $22.58\%$ and $31.25\%$ for each of the RTidy and RBase, respectively. However, the performance of CodeBERT and UniXCoder drop slightly under multi-lingual fine-tuning, but the p-value still shows that the change is significant since it is less than 0.05. For RCombine, the performance is dropped slightly in multi-lingual setting for all models, suggesting that combining the two R styles for method name prediction is not helpful, though separate styles would benefit in multi-lingual setting.
For RCombine, RTidy, and RBase, all p-values are less than $0.05$ (marked with an asterisk), indicating statistically significant differences compared to the monolingual setting, with one exception. The exception is UniXCoder’s performance on RTidy, where the p-value of $0.227$ suggests no statistically significant difference. Notably, GraphCodeBERT consistently shows the best performance across all R datasets, with F1-scores of $0.28$, $0.38$, and $0.41$ for RCombine, RTidy, and RBase respectively, outperforming all other languages. These results are all statistically significant ($p<0.05$).
Overall, the comparison between multilingual training and monolingual fine-tuning reveals that while multilingual training can offer some cross-linguistic generalization, especially for models like GraphCodeBERT, monolingual fine-tuning consistently yields higher performance. The notable differences in performance across different languages and models also highlight the varying degrees to which each model can generalize across languages, with CodeT5, for example, showing the largest gap between multilingual and monolingual settings, struggling more with cross-linguistic generalization compared to others.
CodeT5, built on the T5 architecture, frames code-related tasks as text-to-text problems. This approach allows for a unified model capable of handling various software engineering tasks, including both code summarization and method name prediction. Our results show that each model’s architecture influences its performance differently across tasks. For both tasks, we found that CodeT5’s specific architecture is experiencing significant drop in performance with drop over $50$% and p-value less than $0.001$. This suggests that a unified framework that converts language problem into a text-to-text format will extraordinarily affect the performance of the model under multi-lingual training. It is worth pointing out that CodeT5 also has an early stop mechanism, which will automatically stop fine-tuning if the training results do not improve in relation to a specific round to reduce the fine-tuning time. When fine-tuning under multi-lingual setup, the model takes much longer than mono-lingual fine-tuning, and more epoch to stop. This also suggests that converting the problem to a text-text problem under multi-lingual fine-tuning poses extra challenge for the model to convergence.
4.3. RQ3: Effect of Intra- and Cross-Project Training
4.3.1. Code Summarization
Dataset/Model CodeBERT GraphCodeBERT UniXCoder CodeT5
Cross-project RCombine 4.12 4.28 5.95 5.9
RTidy 3.48 3.51 4.35 5.19
RBase 4.13 3.68 6.58 5.66
Intra-project RCombine 16.56 11.91 19.82 18.36
RTidy 18.78 10.76 18.94 17.64
RBase 31.18 9.46 21.45 16.87
Table 6. Result of code summarization for RQ2 Best results in each row are bold and best results in each column are underlined.
Table 6 reports the results of code summarization using cross-project training across CodeBERT, GraphCodeBERT, UniXCoder, and CodeT5. In cross-project training, data elements from one project are strictly allocated to either training, testing, or validation sets, ensuring that the model is tested on entirely unseen data from different projects.
For the Base R style, UniXCoder outperforms the other models with a BLEU-4 score of $6.58$, followed by CodeT5 with $5.66$, CodeBERT with $4.13$, and GraphCodeBERT with $3.68$. In the Tidy-verse style, CodeT5 again leads with a score of $5.19$, while UniXCoder is close behind at $4.35$. CodeBERT and GraphCodeBERT perform similarly in this style, scoring $3.48$ and $3.51$, respectively. When considering the combined R dataset, which includes both Tidy-verse and Base R, UniXCoder maintains the highest performance with a BLEU-4 score of $5.95$, with CodeT5 closely trailing at $5.9$. GraphCodeBERT slightly outperforms CodeBERT in this combined scenario, scoring $4.28$ compared to $4.12$.
These results demonstrate that UniXCoder mostly delivers superior performance across different R programming styles in a cross-project training scenario, particularly excelling in Base R tasks. CodeT5 also shows competitive results, particularly in the Tidy-verse style. Though UniXCoder and CodeT5 have highest scores in the cross-project setting, their performance has dropped significantly, more than $11$ BLEU-4 scores compared to Table 2. Similar to RQ1, Base R and the combined version have higher scores compared to Tidy-verse. This finding align with the previous findings of intra- vs cross project setting (Zhang et al., 2016a), where intra-project outperforms cross-project setting in most cases.
In contrast, the intra-project results, where the models are tested on the same project they were trained on, demonstrate higher performance. CodeBERT, for example, achieves a significant improvement on the RBase dataset, scoring $31.18$, an improvement from $19.55$ BLEU-4. We observe the highest improvement for CodeBERT followed by CodeT5 compared to other models. Among the two R styles and RCombined, RBase scores increase more than the others.
Overall, the intra-project results underscore the importance of project-specific training, as models tend to perform significantly better when tested on data with a similar distribution. This is particularly evident from the large boosts in the models’ performances from cross-project to intra-project scenarios. GraphCodeBERT, however, while still improving, does not gain as much, suggesting that it may struggle more with the nuances of project-specific contexts. This is aligned with our previous findings in RQ1, where GraphCodeBERT had the lowest score among all models. CodeT5 and UniXCoder, while performing well across both settings, show a slightly smaller margin of improvement, indicating a more consistent but less specialized approach to the task.
4.3.2. Method Name Prediction
Dataset CodeBERT GraphCodeBERT UniXCoder CodeT5
Corss-project RCombine 0.17 0.18 0.19 0.20
RTidy 0.14 0.17 0.15 0.17
RBase 0.16 0.16 0.11 0.19
Intra-project RCombine 0.27 0.27 0.19 0.30
RTidy 0.26 0.27 0.18 0.30
RBase 0.26 0.27 0.20 0.30
Table 7. The result for method name prediction using cross-project training and Intra-project training. Best results in each row are bold and best results in each column are underlined.
Table 7 represents the results for method name prediction in intra- and cross-project setting. The results highlight the consistent trend observed in prior studies involving intra and cross-project setting (Zhang et al., 2016b). Models generally perform better in intra-project settings compared to cross-project settings. This pattern is evident across all four models when applied to the R programming datasets.
In the cross-project training scenario for RCombine, CodeT5 achieved the highest prediction accuracy with a score of $0.20$, followed closely by UniXcoder at $0.19$. CodeBERT and GraphCodeBERT lagged slightly behind, with scores of $0.17$ and $0.18$, respectively.
In the intra-project training context, the performance improved overall, with the best result again coming from CodeT5, which reached $0.30$. Both CodeBERT and GraphCodeBERT scored equally at $0.27$, while UniXcoder recorded a lower score of $0.19$. There is a notable improvement in performance when moving from cross-project to intra-project training for RCombine, with scores increasing by 0.10 for CodeBERT and GraphCodeBERT, and CodeT5. CodeT5 shows the best performance with RCombine in both settings.
In the intra-project setting, where the model is trained and tested on data from the same project, the F1 scores are significantly higher across all datasets and models. For instance, CodeT5 achieves an F1 score of 0.30 on the RBase dataset, compared to 0.19 in the cross-project setting. Similarly, UniXCoder’s performance on the RBase dataset jumps from 0.11 in the cross-project setting to 0.20 in the intra-project setting. This improved performance indicates that models can better capture the idiosyncrasies and specific coding practices within a single project, leading to more accurate predictions.
4.4. RQ4: Code-PLM Performance When Separating R Style Datasets
Figure 2. Performance of baseline models on code summarization and method name prediction tasks across different programming styles in R under monolingual and multilingual setups.
In this section, we examine the performance of four baseline models fine-tuned on three distinct R programming datasets for two tasks of code summarization and name prediction, as shown in Figure 2. The goal is to evaluate how different styles of R and varying training setups impact the models’ ability to accurately summarize code and predict method names. We also examined the statistic significance of the difference between RCombine and two styles RTidy and RBase. Since the population size is not the same across RCombine, RTidy, and RBase we take random samples from RCombine to match the size of other two.
  • For code summarization under monolingual training, the differences in performance are statistic significant for CodeBERT and CodeT5, the differences are only significant between RCombine and RBase for GraphCodeBERT and UniXCoder. When looking at name prediction under monolingual fine-tuning, the difference in performance is only statistic significant between RCombine and RBase for UniXCoder.
  • For code summarization under multi-lingual fine-tuning, the differences in performance are statistic significant for all models except CodeT5, where the performance dropped 5% and 9.4% for RTidy and RBase, respectively. For name prediction under multi-lingual fine-tuning, similar trends were found, where the difference in performance is only statistic significant for one model, which is GraphCodeBERT.
  • For code summarization under cross-project training, the difference in performance is statistic significant between RCombine and RTidy for CodeBERT, GraphCodeBERT and UniXCoder. The differences are more significant under intra-project training, where all of them are statistic significant.
  • For method name prediction using cross-project training, the difference in performance is significant for CodeBERT and UniXCoder. It is only significant for GraphCodeBERT and CodeT5 between RCombine and RBase and RCombine and RTidy, respectively. Unlike code summarization, the differences are less significant under intra-project training, where all of them are not statistically significant different.
In the code summarization task for the monolingual setting, the results indicate that models trained on the Base R dataset outperform those trained on the Tidy-verse and the combined datasets. The Base R models produced the highest scores for all baseline models except GraphCodeBERT, indicating that these models handle Base R syntax more effectively when summarizing code. The improvement in performance for Base R could be a benefit from a syntax structure that is simpler and closer to other programming languages. Tidy-verse-based models, while still performing relatively well, did not match the performance of RBase models, suggesting that the abstract and functional style of Tidy-verse poses additional challenges for the models during fine-tuning phase. The combined dataset, which merges both Base R and Tidy-verse, resulted in the lowest performance. This outcome implies that combining two distinct paradigms introduce complexities that interfere with the model’s ability to generate concise and accurate summaries. The performance shows that the two styles introduce conflicting patterns to the models and make the models struggle to reconcile.
In contrast, in the name prediction task, the performance differences between the datasets are minimal. Whether the models were trained on Base R, Tidy-verse, or the combined dataset, the performance remained relatively similar. Previous work (Ahmed and Devanbu, 2022) shows that identifier plays the most important role in code tasks. This consistency indicates that method name prediction is less sensitive to the specific R style used, but more rely on the identifiers to capture the key information. Developers would use method names across Base R and Tidy-verse that share enough similarities to enable the models to generalize effectively across these paradigms. The combined dataset, although slightly weaker, did not exhibit the performance drop seen in code summarization. The consistency across different R programming styles reinforces the observation that name prediction is less dependent on variations in R programming styles, but mainly dependent on the identifiers.
When shifting to the multilingual setup, the performance for both tasks showed a various degrees of decline. In the code summarization task, the same pattern observed in the monolingual setup persisted: models trained on Base R achieved the highest scores for most baseline models, while the combined dataset performed the worst. Contradicting previous research (Ahmed and Devanbu, 2022), the multilingual environment did not improve the performance of the models, it introduces additional complexity, as the models must handle syntax and structures from multiple programming languages that has different syntax and structure in many ways. This added burden likely impedes their ability to focus on the nuanced differences between Base R and Tidy-verse, making it more difficult to generate accurate code summaries, especially when trained on a mix of both paradigms.
Similarly, in the multilingual name prediction task, the models’ performance remained consistent across datasets, mirroring the monolingual setting. The scores displayed little variation, with Base R, Tidy-verse, and combined datasets all yielding nearly identical results, except for GraphCodeBERT. This further confirms that name prediction is a relatively straightforward task for these models, even when trained in a multilingual setting. It suggests that method names in R are relatively easy for the models to generalize, regardless of the specific programming style or the presence of multiple languages in the training set. Similar to code summarization, the scores are dropped in multilingual settings.
To conclude, this section’s analysis reveals several key trends. First, for code summarization, models trained exclusively on Base R tend to perform better than those trained on Tidy-verse or a combination of both, particularly in a monolingual setup. The Tidy-verse introduces additional complexity, which the models do not handle as effectively, and combining both styles into one dataset exacerbates this issue. In contrast, for method name prediction, the models are largely unaffected by the specific R paradigm, with shows similar performance across all datasets and setups, indicating that this task is less sensitive to the underlying coding style. Finally, while the multilingual setup introduces some performance degradation, particularly in code summarization, the models maintain consistent performance in name prediction, demonstrating their ability to generalize method names across different programming languages and paradigms. These insights highlight the importance of choosing the right dataset and fine-tuning setup for specific tasks, especially in complex environments like R programming.
4.5. RQ5: LLM Results
K=1 K=5 K=10
Models R Python R Python R Python
UniXcoder 0 0 0 0 0 0
codegen-2B-multi 4.32 10.18 6.89 19.51 7.76 26.83
CodeLlama-7B 17.76 25.98 27.74 48.17 33.5 57.93
StarCoder2-7B 19.65 28.9 28.35 51.21 35.4 66.46
Table 8. Pass@k on HumanEval for selected LLMs.
The results for code generation on MultiPL-E (Cassano et al., 2023) and Humaneval (Chen et al., 2021) are shown in Table 8. UniXCoder achieves 0 for both R and Python for all metrics, showing it is not able to produce executable code, in contrary to the LLMs. The performance disparity between R and Python is a key finding of this study. Across all evaluated models, R consistently yields lower pass@k scores compared to Python. For instance, the top-performing StarCoder2-7B model achieves a pass@10 score of $35.4\%$ for R, which, while impressive, falls significantly short of its $66.46\%$ score for Python. This gap is consistently observed across different model sizes and architectures, suggesting a systemic challenge in R code generation. While UniXcoder fails to generate valid solutions for both languages, achieving 0% pass rates across all K values, larger models show more promise but with notable language-dependent differences. CodeLlama-7B and StarCoder2-7B demonstrate superior performance, particularly for Python, with StarCoder2-7B achieving the highest pass rates of 28.9% at K=1 and 66.46% at K=10 for Python code generation.
The improvement in pass@k scores as k increases is particularly noteworthy for R. For StarCoder2-7B, the score nearly doubles from k=1 (19.65%) to k=10 (35.4%). This suggests that allowing multiple generation attempts could significantly enhance the practical utility of these models in R programming environments. It also implies that the models have a broader understanding of potential R solutions but may require multiple attempts to articulate the correct one. The performance trends across model sizes are consistent for both R and Python, with larger models generally performing better. For instance, the jump from codegen-2B-multi to CodeLlama-7B shows a more substantial relative improvement for both R and Python. This performance improvement is expected, as the model has more attempts to create a correct code, which is not specific to R.
4.6. Human evaluation
Model Informativeness Relevance Fluency
CodeBERT 2.83 3.07 3.76
GraphCodeBERT 3.09 3.28 3.80
UniXCoder 3.09 3.20 3.78
CodeT5 3.53 3.59 3.94
Ground truth 3.99 3.97 4.26
Table 9. Human evaluation result for code summarization
Table 9 shows the human evaluation scores received by each model for code summarization task. Each model is evaluated from three aspects: informativeness, relevance, and fluency.
Informativeness: CodeT5 performs the best among the models with a score of $3.53$, which is notably closer to the ground truth ($3.99$) than the other models. This suggests that CodeT5 is more effective at extracting and conveying key information from the code. GraphCodeBERT and UniXCoder tie at $3.09$, indicating they perform similarly in this aspect, while CodeBERT lags slightly behind at $2.83$. The gap between the best-performing model (CodeT5) and the ground truth (0.46) indicates that there is still room for improvement in capturing the full informational content of the code.
Relevance: This aspect assesses how well the generated summaries align with the actual functionality and purpose of the code. Again, CodeT5 leads the pack with a score of $3.59$, closest to the ground truth of $3.97$. GraphCodeBERT follows at $3.28$, with UniXCoder close behind at $3.20$, and CodeBERT at $3.07$. The scores suggest that while all models produce relevant summaries to some degree, CodeT5 is noticeably better at generating summaries that accurately reflect the code’s purpose. The smaller gap between models in this category compared to Informativeness suggests that achieving relevance could be easier than capturing all informative details.
Fluency: This metric evaluates the linguistic quality and readability of the generated summaries. All models perform relatively well in this aspect, with scores clustered more closely together and closer to the ground truth. CodeT5 again leads with $3.94$, close to the ground truth of $4.26$. The other models follow closely: GraphCodeBERT ($3.80$), UniXCoder ($3.78$), and CodeBERT ($3.76$). The high and closely grouped scores suggest that all these models are quite proficient at generating fluent and readable summaries, with less differentiation between them in this aspect compared to Informativeness and Relevance.
These results indicate that CodeT5 consistently outperforms the other models across all three aspects, coming closest to the ground truth in each category. This suggests that CodeT5 has made significant strides in understanding and summarizing R code. The scores also reveal that while these models have achieved impressive fluency in their summaries, there is still a noticeable gap in their ability to fully capture the informative content and relevance of the code compared to human-generated summaries (ground truth).
Model Semantic Similarity
CodeBERT 2.53
GraphCodeBERT 2.66
UniXCoder 2.27
CodeT5 2.83
Ground truth 2.66
Table 10. Human evaluation result for name prediction
Table 10 shows the human evaluation scores for the baseline models. Semantic similarity measures how well the predicted names capture the meaning and purpose of the code elements they are naming. It is a crucial aspect of name prediction, as semantically appropriate names significantly enhance code readability and maintainability. CodeT5 stands out as the top performer with a score of $2.83$, which is notably higher than the ground truth score of $2.66$. This is an intriguing result, suggesting that CodeT5 is capable of generating names that human evaluators find even more semantically appropriate than the original names in some cases. This could indicate that CodeT5 has developed a strong understanding of code semantics and naming conventions, possibly even surpassing human performance in certain contexts. On the other hand, the low score received by the ground truth also indicates that some ground truth answer would make more sense when given the background information of the project.
GraphCodeBERT follows with a score of $2.66$, which interestingly matches the ground truth score exactly. This suggests that GraphCodeBERT’s name predictions are, on average, as semantically similar to the code’s purpose as the original names. CodeBERT comes in third with a score of $2.53$, which is relatively close to the ground truth. While not matching the performance of CodeT5 or GraphCodeBERT, it still demonstrates a decent ability to generate semantically appropriate names. UniXCoder shows the lowest performance among the models with a score of $2.27$, which is noticeably below the ground truth and other models. This suggests that UniXCoder may have more difficulty capturing the semantic nuances required for effective name prediction in this context. The relative performance of these models in name prediction aligns somewhat with their performance in code summarization (from the Table 10), with CodeT5 leading in both tasks. This consistency suggests that the underlying capabilities that make a model good at summarizing code also contribute to its ability to predict semantically appropriate names.
5. Discussions
In this section, we examine several factors influencing the observed performance in our study, particularly in the context of R programming. Key points include the choice of models employed, the impact of dataset, the models’ generalization capabilities, and the effectiveness of different training strategies.
5.1. R as a Low Resource Language
As one of the driving forces behind this early-stage research, we are keenly aware that R, as a low-resource language, faces significant challenges in both training and fine-tuning. For example, in the process of dataset collection, R may not yield as many high-quality training sets as mainstream languages like Python. Also, during the course of this research, we recognize that most current pre-trained models and large language models require vast amounts of text and code snippets for training. Moreover, the specificity and uniqueness of the R language make it difficult for R to appear on a large scale in corpora used for training. For example, on GitHub, a well-known hosting platform, we used the same search criteria to harvest datasets. The results, comparing R with other mainstream languages, are presented in Figure 3.
Figure 3. number of available projects on GitHub. Result is obtained by requesting GitHub through REST API with same query statements except programming language
From Figure 3, JavaScript, Python, and Java have approximately 10 times more GitHub repositories than R. Moreover, we can see that the available R projects is half of the size compare to Ruby, which is considered as low-resource language in the CodeSearchNet (Husain et al., 2019) dataset. The relatively low number of R packages on GitHub compared to other popular programming languages presents significant challenges for training effective language models and LLMs focused on R.
The smaller dataset also poses difficulties for transfer learning and fine-tuning approaches. We can see from Table 4 and Table 5, multi-lingual fine-tuning would not increase the performance even this approach has been showed to be effective in previous studies (Ahmed and Devanbu, 2022).
When adapting general-purpose language models to R-specific tasks, there’s a risk of over-fitting due to the limited examples available. This could hinder the model’s ability to generalize effectively and may lead to sub-optimal performance when dealing with novel R code or projects that integrate multiple programming languages. The reduced diversity in coding styles and patterns within the smaller pool of R packages might also impact the model’s understanding of the broader R ecosystem, including package dependencies and usage patterns.
Listing 3: POV-base
1function(x, filter.dups = FALSE) {
2 checkmate::assertMatrix(x, mode = "numeric", min.cols = 1L, min.rows = 1L)
3
4 df = as.data.frame(t(x))
5
6 colnames(df) = paste0("y", seq_len(ncol(df)))
7
8 if (filter.dups){
9 df = filterDuplicated(df)
10 }
11
12 return(df)
13}
Listing 4: POV-tidy
1function(x, filter.dups = FALSE) {
2 checkmate::assertMatrix(x, mode = "numeric", min.cols = 1L, min.rows = 1L)
3
4 df <- as_tibble(t(x))
5
6 df <- df %>rename_with(~ paste0("y", seq_len(ncol(df))))
7
8 if (filter.dups) {
9 df <- df %>distinct()
10 }
11
12 return(df)
13}
Model Content
CodeBERT Base: Convert a matrix to a data frame
Tidy: matrix to data.frame
GraphCodeBERT Base: Convert a matrix to a data frame
Tidy: matrix to data.frame
UniXCoder Base: lifecycle - experimental options : alt = Experimental
Tidy: rename matrix to df
CodeT5 Base: Converts a numeric matrix to a data.frame
Tidy: rename matrix to df
CodeLlama2 Base: This function takes a numeric matrix as input and returns a data frame
Tidy: This function is used to convert a matrix to a tibble
StarCoder2 Base: transform the data into a data frame with one row per sample and one column per feature
Tidy: The matrix is transposed, so that the rows are the observations and the columns are the variables. The variables are renamed to ‘y1‘, ‘y2‘
Table 11. Summarizations generated by the models
5.2. Impact of Tidy vs. Base
The dichotomy between Tidy-verse and Base R in the R programming presents an interesting challenge for language models and LLMs trained on R code. This division could significantly impact model performance and introduces unique considerations for models working with R. Tidy-verse, a collection of popular R packages designed for data science, introduces a distinct syntax and philosophy that differs from Base R. This split in coding styles and approaches within the R community can affect language models in several ways. Models trained on a dataset that includes code from both R styles, may struggle to maintain consistency in their outputs, potentially mixing idioms. This is more challenging specifically interfere code generation by mixing both paradigms. This could lead to confusion for users expecting pure Tidy-verse or Base R solutions. For example, the summaries generated by CodeLlama2 for both Tidy and Base shows a good understanding of the code’s functionality but the Tidy version shows less information about the function compared to the Base version.
Listings 3 and  4 are the Base and Tidy version of the same function. The function processes a numeric matrix by first validating the input to ensure it meets the minimum criteria. It transposes the matrix, converts it into a data frame, and renames the columns systematically as $y_{1},y_{2},\dots,y_{n}$. Optionally, it filters duplicated rows based on user specification, ensures that the data is structured and ready for further analysis or modeling. The Base R implementation, as shown in the Listing 3 example, follows a more traditional, imperative style, relying on Base R functions like as.data.frame(), t(), and filterDuplicated(). In contrast, the Tidy-verse version, listing 4, embraces a more expressive, declarative approach, utilizing as_tibble(), rename_with(), and distinct() from the Tidy-verse packages.
We provided a list of summarizations from each model in Table 11, where they are asked to summarize a Base and Tidy version of the same function shown in Listings 3 and  4. The pre-trained language models are trained for code summarization and the LLMs are doing this task under a zero-shot setup. The performance of the analyzed models reflects their varying levels of exposure to and familiarity with these R programming idioms. CodeBERT ,GraphCodeBERT and CodeT5, demonstrate a stronger grasp of the underlying concepts when compared to UniXcoder, as evidenced by their precise summaries for both the Base R and Tidy-verse versions. However, the summaries provided by those models are very short in general, missing key information about the function’s underlying task. For example, only CodeT5 in its summary for the Base version, motioned that the matrix need to be numeric. Moreover, the word “rename” is not accurate about the function’s functionality. In contrast, when we look at the summaries provided by the LLMs, one of the main difference is that LLM’s summaries are longer, and contain more information. For instance, CodeLlama provided information about the input and output of the function. Moreover, StarCoder provides summaries about what happened during the execution. On the other hand, StarCoder did not mention that the matrix is transposed into a dataframe when summarizing Tidy version, and missed the numeric part for Base version.
Furthermore, all models’ ability to capture project-level context and constraints, such as domain-specific requirements, implicit dependencies, and established coding conventions, appears to be an area for potential improvement. All of the models missed the part about checkmate::assertMatrix, where it is a function from a larger context. The summaries generated by the models, while accurate in conveying the general purpose of the functions, lack the nuanced understanding of how these code snippets would integrate within the larger context of a real-world R project. This is an existing challenge for other programming languages, as shown in previous studies in the existance of unfamiliar libraries or private repositories (Zhou et al., 2022; Jimenez et al., 2023). Previous work (Yun et al., 2024) also shows that including context information would increase the performance for code summarization task.
Moreover, from the examples provided in listing 3 and listing 4, we can see that the prevalence of Tidy-verse in modern R programming might skew model performance. If the training data heavily favors Tidy-verse code, the model may underperform when dealing with Base R tasks, and vice versa. This imbalance could result in biased code suggestions or completions, potentially steering users towards one paradigm over the other unintentionally, especially when Tidy-verse is heavily used in some domain while barely used in others. The model’s ability to understand and generate idiomatic R code could vary significantly depending on whether it is working with Tidy-verse or Base R constructs.
Finally, building upon our empirical investigation, we uncover an intriguing relationship between data volume and model performance across R syntax paradigms. Despite Tidy R syntax having approximately three times more training data than Base R, our results indicate that this substantial data advantage does not translate into consistently superior model performance. Analysis across multiple models reveals that Base R syntax often achieves comparable or marginally better results: UniXCoder and CodeBERT demonstrate slightly higher performance with Base R (20.93% and 19.55% respectively) compared to Tidy R (18.93% and 17.92%). Only GraphCodeBERT shows a minor advantage when processing Tidy syntax (10.38% vs. 9.56% for Base). Notably, the combined syntax approach generally underperforms compared to either individual syntax, suggesting increased complexity when models must handle both paradigms simultaneously. These findings challenge the assumption that larger training datasets inherently lead to better performance, indicating that the relationship between syntax paradigms and model effectiveness is more nuanced. Factors such as syntax complexity, consistency, or the model’s architectural bias may play more significant roles than data quantity alone in determining performance outcomes, emphasizing that the quality of the data matters, we discussed earlier.
5.3. Intra- and Cross-Project
Analyzing the effects of intra- and cross-project contexts on model’s performance for R language is complex and nuanced. This aspect of model training presents unique challenges and opportunities that can significantly impact the effectiveness of pre-trained models and LLMs working with R.
Cross-project context presents an greater challenge. R is used across various domains, from bioinformatics to finance, each with its own specialized libraries, data structures, and best practices. A model trained on a broad spectrum of R projects might produce generalized code that fails to account for domain-specific requirements or optimizations. For example, code generated for a bioinformatics project might not inherently understand the need for efficient large-scale data processing that is crucial in genomics research. Another significant challenge lies in understanding and maintaining project-specific logic and business rules. Many R projects in professional settings implement complex analytical pipelines or decision-making processes that are unique to their use case. These often rely on domain knowledge that is not explicitly coded but is crucial for correct implementation.
6. Implications and Future Directions
The findings of this study have several implications for R and scientific software developers and researchers.
First, the insights of our research contribute to the methodological improvement of utilizing Code-PLMs in R. We have demonstrated the performance of Code-PLMs when integrating these models into the R ecosystem. As discussed, one of the issues is the lack of training datasets for R. Researchers and developers working on language models for R could employ several strategies to address this issue. Augmenting the GitHub dataset with high-quality code from other reputable sources, such as CRAN or Bioconductor, could significantly expand the training data. Additionally, applying data augmentation techniques could artificially increase the dataset’s size and diversity. Leveraging transfer learning from models trained on larger programming language datasets, combined with careful domain-specific fine-tuning, could help bridge the gap caused by the limited R-specific data. These approaches, while not eliminating the underlying data scarcity issue, can help improve model performance and ensure more robust and accurate language models for R programming.
Furthermore, researchers and practitioners can devise strategies to enhance model performance, specifically with the lack of huge datasets. This could also include optimizing the selection process and adaptation of Code-PLMs or LLMs for different tasks in R. As observed in RQ5, the LLMs performance on code generation for R is falling behind Python. Specific techniques could be developed to generate acceptable code in R. The code generation and creating a benchmark in Tidy-verse style is also another area of research.
Second, our study provides valuable guidance for R language users who seek to leverage the current Code-PLMs. The comparison and evaluation of different pre-trained models in the R environment shed light on their relative strengths and weaknesses. This information empowers practitioners to make informed decisions when selecting the most suitable model for their specific use cases, considering factors such as model accuracy and compatibility with R programming styles. The practitioner and resercher should be aware of the limitations and strengths of the models when intend to use them for R.
As discussed earlier, the Base and Tidy styles have different syntax and combining them in the training dataset could lower the performance. Careful curation of training data to ensure balanced representation of both Tidy-verse and Base R code is crucial, even separating them for specific tasks. Additionally, models might benefit from explicit tagging or context-awareness to distinguish between Tidy-verse and Base R code segments. This could enable more accurate and context-appropriate code generation and completion. Ultimately, the goal would be to develop models that can seamlessly work with both paradigms, understanding their strengths and appropriately suggesting or generating code that aligns with the user’s preferred style or the specific requirements of the task at hand.
Another implication for researchers is exploring ways to develop datasets and techniques to improve the performance of models for project-specific or R specific domains. One potential solution is to develop models that can intake and process broader project context, e.g., by analyzing entire codebases or project documentation before making suggestions or generating code. This could help the model align its outputs more closely with project-specific conventions and requirements. Another approach might involve creating domain-specific models or fine-tuning general R models for particular fields of application. This could help in generating more relevant and optimized code for specific types of R projects, whether they are in finance, healthcare, or scientific research.
The integration of pre-trained models in R opens up opportunities for their application for other low-resource (i.e., the amount of available data is limited) programming languages, and software engineering tasks. Researchers and practitioners working on these topics can draw upon our findings to explore and exploit the potential of pre-trained models within the R ecosystem, leading to improved solutions and advancements in their respective areas, for other code intelligent tasks and other under-studied programming languages, including scientific software.
7. Threats to Validity
External Validity: We only conduct experiments in R and for code summarization and method name prediction, and for code generation using LLMs only. Limiting the number of tasks is mainly due to reducing the bias of building high-quality data for other tasks. Though the approach can be utilized for other code-related tasks in R, we cannot expand the conclusions to other programming languages or tasks. However, one can benefit greatly from the strategy we introduced for separating the two R styles, namely separating the Base R and Tidy-verse.
Internal validity: The internal validity in our work could be related to the dataset collection process and pre-processing steps. In our study, the quality of data is a significant concern. The code and associated documentation within R packages can vary widely in quality. Some packages may be meticulously maintained with high-quality code and comprehensive documentation, providing valuable insights for model training. However, others may be less well-maintained, with misleading, inaccurate, or sparse documentation. This inconsistency can lead to a model that is trained on a mixture of high and low-quality data, potentially reducing its effectiveness and accuracy. To mitigate this issue, it is crucial to implement robust data cleaning and preprocessing steps, possibly including manual review and curation of the dataset. To achieve this, we consulted constantly with an R expert, who has more than ten years of research and development experience in R, ensuring the quality of the collected dataset and the soundness of the approach, specifically the decision on Base or Tidy-verse R repositories. We also manually evaluated the collected data at different steps, to ensure our developed process leads to the desired data, e.g., the code and comments are tokenized correctly. For the tokenization part, we used several tools, and each failed in one part. Finally, we found a reliable grammar rule for R that could be used with a Tree-sitter. The obtained dataset was investigated by the first author and another student, ensuring that the code and comments were separated correctly. These steps were taken in addition to following the advice of the R expert and previously published works to select the inclusion and exclusion criteria to scrape the related GitHub repositories.
There is an uneven spread between Base R and Tidy-verse packages in the dataset. This could lead to a model that is biased towards the coding style and conventions of one over the other, limiting its effectiveness across the diverse R ecosystem. Additionally, as mentioned earlier, the syntax of the two styles is different, which can affect the performance of the models. Therefore, we added RQ4 to evaluate the effect of separating the data related to each programming style. To avoid any issues with the fair comparison of R styles, we conducted manual evaluations and used the same number of records for RBase and RTidy.
Another internal threat is related to selecting and training the models. We followed the recommendation of a recent paper (Niu et al., 2023) to select the Code-PLMs for code summarization and method name prediction. The author who trained the models has extensive experiments with running deep learning models. Therefore, we anticipate minimal threat related to this point.
Construct validity: We use the commonly used evaluation metrics for code summarization and method name prediction. For code generation we also used a benchmark dataset and the Pass@K metrics, which are well studied in the related literature. These metrics are well known in NLP and SE research and therefore we do not anticipate such validity threats.
Another threat could be related to the multilingual training of the models, investigated in RQ2. Although the difference in the distribution of the training data could affect the results, previous research has shown the benefits of multi-lingual training for code summarization and method name prediction (Ahmed and Devanbu, 2022). We also compare the results of multilingual and monolingual fine-tuning, therefore, anticipating low threats related to this setting.
8. Conclusion and Future Studies
In this study, we explored the capabilities of the Code-PLMs for the two tasks of code summarization and method name prediction for R. This is the first attempt towards code intelligence in R for the mentioned tasks. We also examined LLMs for code generation task for R programming language. Both LLMs and Code-PLMs consistently experience varying degrees of performance degradation when processing R code compared to other programming languages. The challenge stems in different syntax of the R styles as well as its project-specific and package-based nature. Interestingly, increasing the dataset size or adding other programming languages for multi-lingual fine-tuning not necessarily helps in improving the performance.
The significance of this study extends far beyond its immediate findings about R language processing. As software development increasingly relies on AI-assisted programming tools, our research addresses a critical gap in understanding how current state-of-the-art language models handle diverse programming languages and paradigms. The limitations we uncovered in processing R code have broader implications for the development of code understanding and generation models that can handle different programming languages. As organizations increasingly adopt AI tools for software development, our findings underscore the importance of carefully evaluating these tools’ capabilities across different programming languages and contexts. The limitations we have identified in R processing could potentially extend to other specialized or less common programming languages, making our study an important reference point for future research in this area.
Acknowledgment
This research is support by a grant from Natural Sciences and Engineering Research Council of Canada RGPIN-2019-05175.
We also sincerely appreciate the help of Dr. Melina Vidoni, who has more than ten years of research and development in R, for her guidance and feedback to curate the R dataset.